OpenStudy (anonymous):

MEDAL AND FAN WILL BE AWARDED PLEASE HELP ME. The length of the hypotenuse of a right triangle is 6√3 . what is the perimeter of the triangle? please show work i have more questions like this.

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

Where c = the hypotenuse and a and b are the other two sides.

OpenStudy (superdavesuper):

@nelsonjedi yes ur equation is right but not sure there is enough info to solve for a and b...

OpenStudy (anonymous):

@superdavesuper yea i know thats the equation @nelsonjedi

OpenStudy (superdavesuper):

any more info e.g. like a picture or something else in the question?

OpenStudy (anonymous):

Well if a^2 + B^2 = C^2 and we know C^2. The perimeter should be A+B+C. How about this solution \[2\times(\sqrt{c ^{2}})\]

OpenStudy (anonymous):

The answer is two times the hypotenuse

OpenStudy (anonymous):

Suoerdave does that work?

OpenStudy (superdavesuper):

@nelsonjedi unfortunately no :(

OpenStudy (anonymous):

so what does work? @superdavesuper

OpenStudy (superdavesuper):

Sorry - no way to solve if there is no other info like a drawing or other details.

OpenStudy (anonymous):

there isnt thats all it says. there has to be some way to solve.

OpenStudy (superdavesuper):

i dont really think so but u can try asking for help from someone more "senior" than me :) good luck!